OpenStudy (anonymous):

Conditional Probability.. I think Suppose 20 numbered balls are in an urn, each with a unique number, 1 through 20 on it. If you draw 4 (and order doesn't matter), approximate the probability that the set of 4 will contain at least 2 consecutive numbers. (for example, k and k+1 for 9 and 10.) Any suggestions on how to set a problem like this up?

OpenStudy (anonymous):

i wonder why it says "approximate"?

OpenStudy (anonymous):

i can think of a (not necessarily right) procedure that might be an approximation, but it would certainly not be exact

OpenStudy (anonymous):

no, i see it is wrong. damn

OpenStudy (anonymous):

lets start anyway, maybe @Zarkon will correct it pick a number assuming it is not one or 20 when you pick the next number there are 2 out of the remaining 19 that could be adjacent if so you are done and the probability is \(\frac{2}{19}\) if not, with probabilty \(\frac{17}{19}\) you pick one that is not adjacent. then the probability that the next one is adjacent to one of the first two is either \(\frac{3}{18}\) or \(\frac{4}{18}\) depending on the first two damn this won't work

OpenStudy (anonymous):

maybe we can do it by passing to the compliment what is the probability that none are consecutive

OpenStudy (zarkon):

\[1-\frac{{20-4+1\choose 4}}{{20\choose 4}}\]

OpenStudy (zarkon):

@satellite73 That is the way to do it number the 4 balls to be chosen in order \[N_1,N_2,N_3,N_4\] now they need to be chosen so that there is at least one ball between them...mark that ball with a star \[N_1\star,N_2\star,N_3\star,N_4\] so we still have 4 choices but treat \(N_1\star,N_2\star,N_3\star\) as 3 balls (not 6). that gives us a total of 17 balls to choose from (when picking 4)...hence \[{ 17\choose 4}={20-4+1\choose 4}\] the denominator is then obviously \(\displaystyle {20\choose 4}\)

OpenStudy (anonymous):

this is why i should stay up later. missed this answer last night